Circular No. 6545

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970111 D. A. Frail, National Radio Astronomy Observatory; S. R. Kulkarni,

California Institute of Technology; L. Nicastro, D. Dal Fiume, M. Orlandini, E. Palazzi, and G. Pizzichini, ITESRE, Bologna; F. Frontera and G. Zavattini, Universitá di Ferrara; L. Amati, M. N. Cinti, E. Costa, M. Feroci, and L. Piro, IAS, Frascati; and R. Jager and J. Heise, SRON, report: “We have detected a single unresolved (at a level of 1′′) radio source coincident with one of the two x-ray sources (source ‘a’) seen toward GRB 970111 by BeppoSAX (IAUC 6539). Observations at 20 cm with the Very Large Array on Jan. 13.58 UT yield the position α = 15^{h}28^{m}44^{s}.57 ± 0^{s}.01, δ = +19^{o}44′35′′.0 ± 0′′.2 (equinox J2000.0) and a flux density of 2.0 mJy. There is no radio source in the error circle of BeppoSAX source ‘b’ above a 3σ level of 0.18 mJy. A radio monitoring program is underway of the entire 10′-radius error circle; a preliminary analysis of 20-cm data taken on Jan. 12.58, 13.58, and 15.57 shows no significant variability for any of the radio sources detected in this larger field at a level of 1 mJy or more.”

K. Hurley, Space Sciences Laboratory, University of California, Berkeley, on behalf of the Ulysses Gamma-Ray Burst Team; and C. Kouveliotou, G. Fishman, and C. Meegan, Marshall Space Flight Center, NASA, on behalf of the Compton Gamma Ray Observatory BATSE Team, report: “A preliminary triangulation of this event has been done using Ulysses and BATSE data. The resulting position is an annulus whose half-width is 0^{o}.045 centered at α = 11^{h}50^{m}56^{s}, δ = +33^{o}21′.8 (equinox 2000.0), with central radius 49^{o}.979. Combining this with the BeppoSAX error circle (IAUC 6533) gives an approximately trapezoidal error box with corners at α, δ = 15^{h}28^{m}20^{s}, +19^{o}30′.2; 15^{h}28^{m}57^{s}, +19^{o}46′.8; 15^{h}27^{m}59^{s}, +19^{o}32′.2; 15^{h}28^{m}38^{s}, +19^{o}49′.7 (equinox 2000.0). Only one of the BeppoSAX sources given on IAUC 6539, source ‘a’, lies within this error box. Further analysis is expected to reduce the width of this error box substantially.”

EG CANCRI Visual magnitude estimates by P. Schmeer, Bischmisheim, Germany

(cf. IAUC 6517): 1996 Dec. 12.91 UT, 12.6; 15.10, 13.7; 15.21, 13.8; 1997 Jan. 16.14, [13.7; 16.83, 13.0:; 17.14, 13.3.

1997 January 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green